OpenStudy (anonymous):

Could someone help me on this problem, I'm a bit lost on it. I know how to do problems like these, but I'm still not understanding what my answer should be. Question: Let f(x) = x + 8 and g(x) = x2 − 6x − 7. Find f(g(2)). Answer Choices: A) −7 B) -3 C) 10 D) 33

OpenStudy (anonymous):

@Nnesha @nincompoop @dan815

OpenStudy (maddy1251):

Do you know how to get f(g(x))?

OpenStudy (maddy1251):

Or to get g(f(x))

OpenStudy (anonymous):

Yes, you would have to plug in the 2 for both f(x) and g(x). For both equations.

OpenStudy (maddy1251):

Alright, good. I just learned these a month ago, so I am no pro, but I will try to help you the best I can. :-) I am fairly good with these.

OpenStudy (anonymous):

Would you like me to show you my work and then you could correct me on any mistakes I made?

OpenStudy (maddy1251):

Question: Let f(x) = x + 8 and g(x) = x2 − 6x − 7. Find f(g(2)). Answer Choices: A) −7 B) -3 C) 10 D) 33 Okay, so it is saying f(g(2)), which is like f(g(x)), so we should probably try to find that first.

OpenStudy (maddy1251):

If you want, sure :)

OpenStudy (anonymous):

Alright haha. So, we'll start with g(x) first. You'd plug in the 2 for all of the x variables. g(2) = (2)^2 - 6 (2) -7 You'd then simplify: g(2) = 4 - 12 - 7 You'd then simplify again: g(2) = -15 ??

OpenStudy (anonymous):

I see I'm having trouble with it if it's not even an answer choice lol.

OpenStudy (maddy1251):

Well, I could see where you are going with. But thats not how we solve for these problems, because they want f of g of x. So, we have to combine f(x) and g(x), THEN we can substitute what f(g(2)) is :-)

OpenStudy (maddy1251):

Question: Let f(x) = x + 8 and g(x) = x2 − 6x − 7. Find f(g(2)). we have x+8 and x^2-6x-7. we have to multiply the polynomials together in order to get f(g(x)).

OpenStudy (maddy1251):

(It seems confusing at first, I used to do what you are doing!)

OpenStudy (maddy1251):

So f(g(x)) will be x+8(x^2-6x-7))

OpenStudy (maddy1251):

Let me multiply these real quick

OpenStudy (anonymous):

Okay, also, you'd substitute the variables for 2 correct? I want to make sure that's correct at least.

OpenStudy (maddy1251):

We will not substitute in f(g(2)) until we combine the polynomials. But you will substutute that in :)

OpenStudy (maddy1251):

First, we need to multiply them together to make it just one polynomial instead of having f(x) and g(x), cause it wants f(g(x))

OpenStudy (maddy1251):

I am combining them now :)

OpenStudy (anonymous):

Oh okay, so how would we do that?

OpenStudy (anonymous):

Oh okay, take your time.

OpenStudy (maddy1251):

I only have one hand to do this cause I injured my other but I will help you, no worries :) Okay, so we have the two polynomials of x^2-6x-7 and x+8

OpenStudy (maddy1251):

What we want to do is multiply them together

OpenStudy (maddy1251):

This will give us the polynomial we need to solve for f(g(2))

OpenStudy (maddy1251):

Because it starts with f(g(x)), with is f(x) x g(x), basically. Make sense?

OpenStudy (anonymous):

Yes, they're both just combined together correct?

OpenStudy (maddy1251):

If it was just g(x), we would just solve for g(2), but it wants f(g(2)).

OpenStudy (maddy1251):

Through multiplicaton, yes.

OpenStudy (maddy1251):

After we combine them, we should get x^3-6x^2-7x-56, if my calculations are correct.

OpenStudy (maddy1251):

This will be our f(g(x))

OpenStudy (maddy1251):

So f(g(x))=x^3-6x^2-7x-56. NOW, try plugging in the 2.

OpenStudy (anonymous):

Could you actually tell me how you combined them together before we actually get into the problem? I want to make sure I do it correctly next time haha. :) If you don't mind.

OpenStudy (maddy1251):

Certainly! Glad you want to learn.

OpenStudy (maddy1251):

First, set up x^2-6x-7 and then multiply it by x+8. We will start multiplying x^2-6x-7 by 8, just like we would regurlar multiplication. That will give us 8x^2-48x-56 (I actually just caught myself on a mistake, glad you wanted me to show you!) THEN we will mulitply by the 'x' term of x+8. Put a 0 under our 56 in the above term, multiply out by the same function. We will get x^3-6x^2-7x. NOW lets combine tem (Which I didn't do, oh goodness.) Drop everything down, we will get x^3-6x^2-41x-56. This should be the correct formula now.

OpenStudy (maddy1251):

Wait, x^3-2x^2-41x-56. My bad.

OpenStudy (maddy1251):

Let me go over my stuff, just to make sure I am right. Before I substitute in.

OpenStudy (maddy1251):

Cause I think I went wrong somewhere, I confused myself with not adding my terms. Lol I am sorry!

OpenStudy (anonymous):

Hmm, I'm still lost on that, I'm sorry. So for the first equation x^2-6x-7, what do you mean by when you multiplied it by x + 8? How would you do that with another equation? I wrote down your steps by the way.

OpenStudy (anonymous):

It's okay lol, I'm glad you're trying to help me so we both can understand lol.

OpenStudy (maddy1251):

Say you have f(x)= x-10 and g(x)=x^2+5x-2. We would multiply them together, so take x^2+5x-2 times x-10. which would yield x^3-15x^2-48x+20. That would be f(g(x)

OpenStudy (maddy1251):

Okay, I know where I WENT WRONG haha it just dawned on me. Calculus fried my brain.

OpenStudy (maddy1251):

We want to find f(g(x)). WELL in order to do that, pretend that g(x is a number.

OpenStudy (maddy1251):

When we have a number for 'x', we just substitute it in for where ever 'x' is.

OpenStudy (maddy1251):

So when we want to find f(g(x)), we just put g(x) into 'x' for f(x)

OpenStudy (maddy1251):

My apologizes for confusing you any, however applying and combining polynomials through multiplication is a good concept that should always be learned.

OpenStudy (maddy1251):

lets go back to our equations we have.

OpenStudy (maddy1251):

No wonder why my answer was completely off. Lol I will get you the answer now.

OpenStudy (maddy1251):

Question: Let f(x) = x + 8 and g(x) = x2 − 6x − 7. Find f(g(2)). soo for f(g(x)) our 'x' for f(x) should be g(x). Right? =x+8 then we will have x^2-6x-7+8 because we are substituing in g(x) for 'x' in f(x).

OpenStudy (maddy1251):

x^2-6x+1 is now our f(g(x))

OpenStudy (maddy1251):

NOW we can solve for f(g(2)). x^2-6x+1 = 2^2-6(2)-1 = -7.

OpenStudy (maddy1251):

Question: Let f(x) = x + 8 and g(x) = x2 − 6x − 7. Find f(g(2)). Answer Choices: A) −7 <----- This is the answer B) -3 C) 10 D) 33

OpenStudy (maddy1251):

I was thinking of something else and got too caught up in what I said. I am going to delete what I said because its irrelevant to this question particularly.

OpenStudy (maddy1251):

Or I'll leave them, since you wanted notes. Anyways, when given any function of f(g(x)) or g(f(x)), it is basically saying to substitute in the equation into for 'x'. We could try reversing to see if you understand?

OpenStudy (anonymous):

Sorry, I'm reading everything over so I understand, just a moment lol.

OpenStudy (maddy1251):

You're fine. I'm sorry if I confused you any, I made a mistake in my understanding and started doing another concept, when all we had to do was substitute in g(x) for 'x' in f(x).

OpenStudy (anonymous):

Oh yes, let's try the reverse to see if I understand it more as well lol.

OpenStudy (maddy1251):

Okay this time I won't mess us. Promise. Do you still want to use 2 as our subsituting variable?

OpenStudy (maddy1251):

*up

OpenStudy (anonymous):

Well, wouldn't we use 2 because that's the number given? What other number could we use?

OpenStudy (maddy1251):

You could use any number, unless they give you a number like they did with f(g(2)). We will still use 2. Lets state the functions Question: Let f(x) = x + 8 and g(x) = x2 − 6x − 7. Find f(g(2)). f(x)=x+8 g(x)=x^2-6x-7. g(f(x)) We will put in f(x) for the variable 'x' in g(x) ; meaning where it shows x squared, x+8 will go there. Remember, though, that we will perform whatever function to the variable being put in, as if it was a whole number such as 2.

OpenStudy (maddy1251):

Now g(f(x)) will be g(x+8)= (x+8)^2-6(x+8)-7. See?

OpenStudy (anonymous):

Oh, because that's f(x) polynomial?

OpenStudy (anonymous):

So what you're doing it sticking f(x)'s polynomial into the g(x) one. I was so confused. xD No I understood what you were doing.

OpenStudy (maddy1251):

Exactly, and for g(f(x)), we use x+8 as the variable. Like how you were putting in '2' for 'x' for g(x), before we put in '2' for g(f(x)), we need to find g(x+8), to get g(f(x)).

OpenStudy (maddy1251):

Yes. I should have wrote it like I just did. My bad

OpenStudy (maddy1251):

When we had f(g(x)), we had f(x^2-6x-7) into our x+8.

OpenStudy (maddy1251):

x= x^2-6x-7. Then we have x^2-6x-7+8.

OpenStudy (maddy1251):

Combine any like terms, which the only ones is the -7 and 8, which when added give us 1.

OpenStudy (anonymous):

Why wouldn't you add the x as well? When you added on just the + 8?

OpenStudy (maddy1251):

Then we got x^2-6x+1.

OpenStudy (maddy1251):

Because that 'x' becomes x^2-6x-7. It's like a number. Lets say g(x) just equaled 7. We wanted to find f(g(x)). So, f(7)= 7+8, we don't add or multiply the 'x' to the 7, do we? No. We are just substituing in. That is all. However, if that 'x' was squared, we would of had to of squared the polynomial. Just like if it was x^2+8, and f(g(x))= f(7). We would then have 7^2+8, because that is the nature of 'x'.

OpenStudy (maddy1251):

Its just like having a numerical variable. Just substitute it in for 'x', you don't need to add or do anything, unless it is squared or possbily cubed, which I haven't encountered cubing yet.

OpenStudy (maddy1251):

Then you would just square it, and then go on with your calcualtions.

OpenStudy (maddy1251):

But that isn't the case for this, so we don't need to worry about it. :-)

OpenStudy (maddy1251):

I hope that helps you out and makes sense. That's the way my Calculus professor taught us.

OpenStudy (anonymous):

Sorry.

OpenStudy (anonymous):

My sister needed to do something on the computer, I'm so glad you're not gone yet lol.

OpenStudy (maddy1251):

It's okay! Do you have other problems or any other questions? I love helping with these.

OpenStudy (anonymous):

Let me read what you had said before lol, just a second. Then we can go over another one. :)

OpenStudy (maddy1251):

Alright. I gave you a medal for working with me, even though I was wrong at first. But, that is the joy of math. :-)

OpenStudy (anonymous):

Lol, I will give you one as well! You've been a great help. One last question, then I'll close this one and tag you in another one. :) So, this would be the final equation: x^2 - 6x - 7 + 8 ?? What would I do next, because I now know what you did. Do I substitute 2 into the variables and then it should come out to -7 ??

OpenStudy (maddy1251):

Yep! You can plug it in and see for yourself. We now know what f(g(x)) equals, so we can find what it equals with '2'.

OpenStudy (anonymous):

Omggg, I see it lol. I did the calculations. I feel so great about the problem now. I was very confused!

OpenStudy (maddy1251):

Because then it becomes f(g(2))=x^2-6x-7+8. With 2^2-6(2)-7+8.

OpenStudy (maddy1251):

Isn't it the best feeling figuring it out? :)

OpenStudy (anonymous):

Yes, instead of just asking for the answer. xD Anyways, I'll close this problem, then we'll go over another one. :) Thank you for helping me with this one!

OpenStudy (maddy1251):

Okay, perfect. Not a problem, glad to help :-)